The z-value used in a 80% Confidence Interval is

1.645

2.575

1.282

1.96
Answer #1

Solution :

Given that,  

At 80% confidence level the z is ,

\alpha = 1 - 80% = 1 - 0.80 = 0.20

\alpha / 2 = 0.20 / 2 = 0.10

Z\alpha/2 = Z 0.10 = 1.282

z-value = 1.282